# Characterizing the support of semiclassical measures for higher-dimensional cat maps

### v1

@@ 4.1.3. Preliminary bound on $B_{\mathrm{w}}$
Note that $\psi_{k}\in S_{L,0,0}(\mathbb{R}^{2n})$ uniformly in $k$ for any coisotropic $L\subset\mathbb{R}^{2n}$. Now we use $\psi_{k}$ to split up $\operatorname{Op}_{h}(\tilde{b}_{-})\operatorname{Op}_{h}(\tilde{b}_{+})$, setting
$P_{k}\coloneqq\operatorname{Op}_{h}(\tilde{b}_{-})\operatorname{Op}_{h}(\psi_{k}^{2})\operatorname{Op}_{h}(\tilde{b}_{+}).$
@@
(4.18) $\operatorname{supp}\tilde{b}_{-}\psi_{k}\subset B_{\frac{r\sqrt{n}}{2}}\left(\frac{rk}{2}\right)\cap\left(\bigcap_{l=0}^{T_{1}-1}A^{-l}\operatorname{supp}\tilde{b}_{z_{l}^{-}}\right).$
@@
$\pi_{+}\left(\operatorname{supp}\left(\tilde{b}_{-}\psi_{k}\right)\right)\subset B_{\frac{Pr\sqrt{n}}{2}}\left(\pi_{+}\left(\frac{rk}{2}\right)\right)\cap\left(\bigcap_{l=0}^{T_{1}-1}A^{-l}\tilde{\pi}_{+}\left(\operatorname{supp}\tilde{\varphi}_{z_{l}^{-}}\right)\right).$
2. Let $\omega_{1},\omega_{2}\in\pi_{+}(\operatorname{supp}(\tilde{b}_{-}\psi_{k}))$. We show by induction that $|A^{l}\omega_{1}-A^{l}\omega_{2}|\leq Pr\sqrt{n}$ for each $10\leq l\leq T_{1}-$ 1. Clearly this holds for $l=0$. Now suppose $|A^{l}\omega_{1}-A^{l}\omega_{2}|\leq Pr\sqrt{n}$ for some $10\leq l<T_{1}-$ 1. We know that $A^{l+1}\omega_{1},A^{l+1}\omega_{2}\in\tilde{\pi}_{+}(\operatorname{supp}\tilde{\varphi}_{z_{l+1}^{-}})$. From (4.19), there exists a $\vec{j}\in(W^{\perp\sigma}+\mathbb{Z}^{2n})\cap W_{+}$ such that
@@
From (4.3), there exists some $C_{0}>0$ such that for all $10\leq l\leq T-$ 1, we have $|\omega_{1}-\omega_{2}|\leq C_{0}\Lambda^{-l}Pr\sqrt{n}$. Importantly,
@@
for $\delta^{\prime}=\min\left\{1-\rho-2\varepsilon,d\left(\rho-1/2-\varepsilon\right)\right\}$. By our choice of $11/2<\rho<$ 1, for $\varepsilon$ sufficiently small, $\delta^{\prime}>0$.
@@
Let $F_{+}$ be the sum of generalized eigenspaces of $A$ corresponding to $\lambda_{1},\ldots,\lambda_{l}$, $F_{-}$ the sum of eigenspaces corresponding to $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$, and $F_{0}$ the sum of generalized eigenspaces of $A$ corresponding to $\lambda_{l+1},\ldots,\lambda_{n},\lambda_{l+1}^{-1},\ldots,\lambda_{n}^{-1}$. Now set $E_{+}$, $E_{-}$, and $E_{0}$ to be the real part of these spaces, i.e.,
@@
Using Gelfand’s formula, for all $\varepsilon>0$,
(5.4) $\|A^{k}\|=\mathcal{O}(\Lambda^{k(1+\varepsilon)}),\quad\|A^{k}|_{L_{-}}\|=\mathcal{O}(\gamma^{k(1+\varepsilon)})\quad\text{as }k\rightarrow\infty.$
As $A^{-1}$ has the same eigenvalues as $A$, we also know
$\|A^{-k}\|=\mathcal{O}(\Lambda^{k(1+\varepsilon)}),\quad\|A^{-k}|_{L_{+}}\|=\mathcal{O}(\gamma^{k(1+\varepsilon)})\quad\text{as }k\rightarrow\infty.$
@@
Suppose $v$ is an eigenvector with eigenvalue $\lambda$ and $u$ is a generalized eigenvector with eigenvalue $\gamma$. If $\sigma(v,u)\neq 0$, then $\gamma=\lambda^{-1}$.
@@
$\sigma(v,u)=\sigma(Av,Au)=\lambda\gamma\sigma(v,u).$
Therefore, $\gamma=\lambda^{-1}$. More generally, suppose $u$ is a generalized eigenvector of rank $\beta+1$. Then there exists generalized eigenvectors $u_{1},\ldots,u_{\beta}$ such that
$(A-\gamma I)u=u_{1},\quad(A-\gamma I)u_{1}=u_{2},\quad\ldots,\quad(A-\gamma I)u_{\beta-1}=u_{\beta},\quad(A-\lambda I)u_{\beta}=0.$
@@
$\sigma(v,u)=\sigma(Av,Au)=\lambda\sigma(v,u_{1})+\lambda\gamma\sigma(v,u).$
If $\lambda\sigma(v,u_{1})=0$, we have $\gamma=\lambda^{-1}$. If not, we repeat the above argument with $\sigma(v,u)$ replaced by $\sigma(v,u_{1})$. We continue to run this argument, stopping and concluding $\gamma=\lambda^{-1}$ if possible. If not, then we eventually conclude $\sigma(u,v_{\beta})\neq 0$. As $v_{\beta}$ is a standard eigenvector, we have $\gamma=\lambda^{-1}$. ∎
Following from the definitions of $E_{\pm}$ and $E_{0}$ and from Lemma 5.1, we see that $E_{\pm}\subset E_{0}^{\perp\sigma}$ and $E_{0}\subset E_{\pm}^{\perp\sigma}$. Thus, $L_{\pm}^{\perp\sigma}\subset L_{\pm}$; in other words, $L_{\pm}$ are coisotropic. Additionally, we see $E_{+}\oplus E_{-}$ is symplectic and $E_{\pm}$ are Lagrangian subspaces of $E_{+}\oplus E_{-}$
@@ 5.1.1. Safe sets
Let $R>\sqrt{2n}$ and let $S_{R}$ be the $(2n-1)$-sphere of radius $R$ centered at the origin. Denote by $\pi:\mathbb{R}^{2n}\to\mathbb{T}^{2n}$ the projection map onto the torus. Define
@@
We give only the proof for $b_{-}$ as the proof for $b_{+}$ follows similarly. By Lemma 2.8, it suffices to show for $i=1,2$ and $10\leq k\leq T_{1}-$ 1, each $S_{L_{-},\rho+\frac{\varepsilon}{2},\rho^{\prime}+\frac{\varepsilon}{2}}$-seminorm of $b_{i}\circ A^{k}$ is bounded uniformly in $k$ and $N$.
@@ 5.2.1. Porosity
As we shift by $z$, it suffices to only consider $I$ centered at the origin. Suppose towards a contradiction that $R$ and $\nu_{0}$ do not exist. Then there exists a sequence of points $\mathrm{z}_{m}\in\mathbb{T}^{2n}$ and lines $I_{m}$ of length $m$ centered at the origin such that for all $\tau\in I_{m}$, $K_{j}$ does not contain $\{\varphi^{t}_{+}(\varphi^{\tau}_{+}(z_{m})):|t|\leq 1/m\}$. We pass to convergent subsequences and assume $z_{m}\rightarrow z_{\infty}$ and $I_{m}$ converges to $\mathbb{R}\tau$ for some $\tau\in\mathbb{R}^{l}$, i.e, the direction of $I_{m}$ converges to the direction of $\mathbb{R}\tau$. Rescale $\tau$ such that for $v_{\infty}=\sum_{i=1}^{l}\tau_{i}v_{i}^{+}$, $|v_{\infty}|=1$. As $K_{j}^{\circ}$ is safe, there exists $c\in\mathbb{R}$ such that $z_{\infty}+cv_{\infty}\mod\mathbb{Z}^{2n}\in K^{\circ}_{j}$. In other words, $\varphi_{+}^{c\tau}(z_{\infty})\in K_{j}^{\circ}$. Thus, for sufficiently large $m$, there exists $\tau_{m}\in I_{m}$ such that $\{\varphi^{t}_{+}(\varphi^{\tau_{m}}_{+}(z_{m})):|t|\leq 1/m\}\subset K_{j}^{\circ}$, a contradiction.
@@
As the right-hand side converges to +$\infty$+ ∞ as $h\rightarrow 0$, by taking sufficiently small $h$, we make the assumption that $0<j<T_{1}$.
@@
For sufficiently small $\varepsilon$, the right-hand-side converges to 00. Using (5.12) and the fact that $K_{w_{j}^{-}}\cap\operatorname{supp}b_{w_{j}^{-}}=\emptyset$, we know $A^{j}(\varphi^{t}_{+}(z)+v)\notin\operatorname{supp}b_{w_{j}^{-}}$ for sufficiently small $h$. From the definition of $b_{-}$, it follows that $\varphi^{t}_{+}(z)+v\notin\operatorname{supp}b_{-}$. ∎
@@ 5.2.2. Application of fractal uncertainty principle
Although $P_{k}$ is defined differently from §4 in this section, the original proof of [dyatlov2021semiclassical]*Lemma 4.5 holds for both definitions. On the other hand, although the statement of Lemma 5.7 is the same as [dyatlov2021semiclassical]*Lemma 4.6, the proof is different. In [dyatlov2021semiclassical], the authors assume a spectral gap condition for $A$. As a result, their lemma follows from an application of the 1-dimensional fractal uncertainty principle. However, as we made no such assumption, our lemma necessitates the higher-dimensional fractal uncertainty principle.
@@ 6.1. Short Periods
[…]
@@
We give a brief summary of their argument. Each summand in (6.7) with $|t|\leq P/4-o(P)$ behaves “locally,” concentrating at the origin in the semiclassical limit. On the other hand, the summands with $P/4+o(P)\leq|t|\leq P/2$ behave “ergodically,” equidistributing on the torus in the semiclassical limit. Finally, the terms close to ±$\pm P/4$ are between local and ergodic, but contribute negligibly in the semiclassical limit. Thus, the sum over $|t|\leq P/4$ weakly converges to $\tfrac{1}{2}\delta(x_{1},\xi_{1})$, while the sum over $P/4<|t|\leq P/2$ weakly converges to $\tfrac{1}{2}dx_{1}d\xi_{1}$.
@@ 6.3. Preliminary calculations
[…]
@@
$\left\langle U_{\omega}G_{h},M_{A}G_{h}\right\rangle_{L^{2}}=\frac{\sqrt{2}}{\sqrt{\operatorname{Tr}(A)}}e^{-\frac{1}{2h\operatorname{Tr}(A)}\left\langle A^{-1}\omega,\omega\right\rangle}e^{\frac{i}{2h\operatorname{Tr}(A)}\left\langle AR\omega,\omega\right\rangle},$
where $R=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$.
@@
[…]
for $R=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$. ∎
@@
There exists $10<c_{0}<$ 1, depending only on $A$, such that for all $l\in\mathbb{Z}^{2}\setminus 0$,
@@
Using the above lemma, we can prove the following. The result generalizes [FNdB]*Proposition 1 to allow for a perturbation $c$.
@@
Suppose $c\in\mathbb{R}^{2}$. Then for all $q\in\mathbb{Z}$,
(6.13) $\sum_{l\in\mathbb{Z}^{2}}\left|\left\langle M^{q}_{A}G_{h},U_{l+c}G_{h}\right\rangle_{L^{2}}\right|\leq C(h\lambda^{|q|/2}+\lambda^{-|q|/2}).$
For $|c|\ll\min(1,\lambda^{-|q|}(h\log h^{-1})^{-\frac{1}{2}})$ and $q\in\mathbb{Z}$ with $h\log h^{-1}\ll\lambda^{-|q|}$,
(6.14) $\sum_{l\in\mathbb{Z}^{2}\setminus 0}\left|\left\langle M^{q}_{A}G_{h},U_{l+c}G_{h}\right\rangle_{L^{2}}\right|\leq Ch\lambda^{|q|/2}.$
Finally, for $|c|\ll 1$ and all $q\in\mathbb{Z}$,
(6.15) $\sum_{l\in\mathbb{Z}^{2}\setminus 0}\left|\left\langle M^{q}_{A}G_{h},U_{l+c}G_{h}\right\rangle_{L^{2}}\right|\leq C(h\lambda^{|q|/2}+(h\log h^{-1})^{\frac{1}{4}}).$
For each inequality, $C$ does not depend on $c$ or $q$.
@@
$\left|\left\langle M_{A}^{q}G_{h},U_{l+c}G_{h}\right\rangle_{L^{2}}\right|=\frac{\sqrt{2}}{\sqrt{\lambda^{q}+\lambda^{-q}}}e^{-\frac{a_{+}(l+c)^{2}\lambda^{-q}+a_{-}(l+c)^{2}\lambda^{q}}{2h(\lambda^{q}+\lambda^{-q})}}.$
We assume that $q\geq 0$. If instead $q<0$, the following argument is the same, up to exchanging $a_{+}(\cdot)$ and $a_{-}(\cdot)$ and replacing $q$ with $-q$.
@@
$\left\{z\in\mathbb{R}^{2}:k-1\leq\frac{a_{+}(z+c)^{2}\lambda^{-q}+a_{-}(z+c)^{2}\lambda^{q}}{2h(\lambda^{q}+\lambda^{-q})}<k\right\}$
$\quad\subset\left\{z\in\mathbb{R}^{2}:\lambda^{-2q}a_{+}(z+c)^{2}+a_{-}(z+c)^{2}<4kh\right\}$
$\quad\subset\left\{z\in\mathbb{R}^{2}:|a_{+}(z+c)|<2\lambda^{q}\sqrt{kh}\text{ and }|a_{-}(z+c)|<2\sqrt{kh}\right\}\eqqcolon S_{k}.$
@@
(6.16) $\sum_{l\in\mathbb{Z}^{2}}\left|\left\langle M^{q}_{A}G_{h},U_{l+c}G_{h}\right\rangle_{L^{2}}\right|\leq C\lambda^{-q/2}\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}.$

### v2

@@ 4.1.3. Preliminary bound on $B_{\mathrm{w}}$
[…]
@@
(4.18) $\operatorname{supp}\tilde{b}_{-}\psi_{k}\subset B_{\frac{r\sqrt{n}}{2}}\left(\frac{rk}{2}\right)\cap\left(\bigcap_{l=0}^{T_{1}-1}A^{-l}\operatorname{supp}\tilde{b}_{\mathfrak{z}_{l}^{-}}\right).$
@@
$\pi_{+}\left(\operatorname{supp}\left(\tilde{b}_{-}\psi_{k}\right)\right)\subset B_{\frac{Pr\sqrt{n}}{2}}\left(\pi_{+}\left(\frac{rk}{2}\right)\right)\cap\left(\bigcap_{l=0}^{T_{1}-1}A^{-l}\tilde{\pi}_{+}\left(\operatorname{supp}\tilde{\varphi}_{\mathfrak{z}_{l}^{-}}\right)\right).$
2. Let $\omega_{1},\omega_{2}\in\pi_{+}(\operatorname{supp}(\tilde{b}_{-}\psi_{k}))$. We show by induction that $|A^{l}\omega_{1}-A^{l}\omega_{2}|\leq Pr\sqrt{n}$ for each $0\leq l\leq T_{1}-1$. Clearly this holds for $l=0$. Now suppose $|A^{l}\omega_{1}-A^{l}\omega_{2}|\leq Pr\sqrt{n}$ for some $0\leq l<T_{1}-1$. We know that $A^{l+1}\omega_{1},A^{l+1}\omega_{2}\in\tilde{\pi}_{+}(\operatorname{supp}\tilde{\varphi}_{\mathfrak{z}_{l+1}^{-}})$. From (4.19), there exists a $\vec{j}\in(W^{\perp\sigma}+\mathbb{Z}^{2n})\cap W_{+}$ such that
@@
From (4.3), there exists some $C_{0}>0$ such that for all $0\leq l\leq T-1$, we have $|\omega_{1}-\omega_{2}|\leq C_{0}\Lambda^{-l}Pr\sqrt{n}$. Importantly,
@@
for $\delta^{\prime}=\min\left\{1-\rho-2\varepsilon,d\left(\rho-1/2-\varepsilon\right)\right\}$. By our choice of $1/2<\rho<1$, for $\varepsilon$ sufficiently small, $\delta^{\prime}>0$.
@@
The differences between the two argument stem from the fact that we do not assume a spectral gap. In  [dyatlov2021semiclassical], the authors use the eigenspaces of the unique largest and smallest eigenvalues, each of which is 1-dimensional. The analogue in our proof is the sum of the generalized eigenspaces of the largest and smallest eigenvalues. As these spaces can be higher-dimensional, we prove multidimensional versions of the results from [dyatlov2021semiclassical]. Our argument picks up some additional small technicalities as we must consider generalized eigenvectors and complex eigenvectors.
@@
Let $F_{+}$ be the sum of generalized eigenspaces of $A$ corresponding to $\lambda_{1},\ldots,\lambda_{l}$, $F_{-}$ the sum of eigenspaces corresponding to $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$, and $F_{0}$ the sum of generalized eigenspaces of $A$ corresponding to $\lambda_{l+1},\ldots,\lambda_{n},\lambda_{l+1}^{-1},\ldots,\lambda_{n}^{-1}$. Now set $_{+}, EE_{-}$, and $E_{0}$ to be the real part of these spaces, i.e.,
@@
Using Gelfand’s formula, for $\varepsilon>0$ and $k$ sufficiently large depending on $\varepsilon$,
(5.4) $\|A^{k}\|\leq\Lambda^{k(1+\varepsilon)},\quad\|A^{k}|_{L_{-}}\|\leq\gamma^{k(1+\varepsilon)}.$
As $A^{-1}$ has the same eigenvalues as $A$, we also know for $k$ sufficiently large
$\|A^{-k}\|\leq\Lambda^{k(1+\varepsilon)},\quad\|A^{-k}|_{L_{+}}\|\leq\gamma^{k(1+\varepsilon)}.$
@@
Suppose $v$ is an eigenvector with eigenvalue $\lambda_{v}$ and $u$ is a generalized eigenvector with eigenvalue $\lambda_{u}$. If $\sigma(v,u)\neq 0$, then $\lambda_{u}=\lambda_{v}^{-1}$.
@@
$\sigma(v,u)=\sigma(Av,Au)=\lambda_{v}\lambda_{u}\sigma(v,u).$
Therefore, $\lambda_{u}=\lambda_{v}^{-1}$. More generally, suppose $u$ is a generalized eigenvector of rank $\beta+1$. Then there exists generalized eigenvectors $u_{1},\ldots,u_{\beta}$ such that
$(A-\lambda_{u}I)u=u_{1},\quad(A-\lambda_{u}I)u_{1}=u_{2},\quad\ldots,\quad(A-\lambda_{u}I)u_{\beta-1}=u_{\beta},\quad(A-\lambda_{v}I)u_{\beta}=0.$
@@
$\sigma(v,u)=\sigma(Av,Au)=\lambda_{v}\sigma(v,u_{1})+\lambda_{v}\lambda_{u}\sigma(v,u).$
If $\lambda_{v}\sigma(v,u_{1})=0$, we have $\lambda_{u}=\lambda_{v}^{-1}$. If not, we repeat the above argument with $\sigma(v,u)$ replaced by $\sigma(v,u_{1})$. We continue to run this argument, stopping and concluding $\lambda_{u}=\lambda_{v}^{-1}$ if possible. If not, then we eventually conclude $\sigma(v,u_{\beta})\neq 0$. As $u_{\beta}$ is a standard eigenvector, we have $\lambda_{u}=\lambda_{v}^{-1}$. ∎
Following from the definitions of $E_{\pm}$ and $E_{0}$ and from Lemma 5.1, we see that $E_{\pm}\subset E_{0}^{\perp\sigma}$ and $E_{0}\subset E_{\pm}^{\perp\sigma}$. Thus, $L_{\pm}^{\perp\sigma}\subset L_{\pm}$; in other words, $L_{\pm}$ are coisotropic. Additionally, we see $E_{+}\oplus E_{-}$ is symplectic and $E_{\pm}$ are Lagrangian subspaces of $E_{+}\oplus E_{-}$ .
@@ 5.1.1. Safe sets
Let $R>\sqrt{2n}$ and let $S_{R}$ be the (2n1)(2n-1)-sphere of radius $R$ centered at the origin. Denote by $\pi:\mathbb{R}^{2n}\to\mathbb{T}^{2n}$ the projection map onto the torus. Define
@@
We give only the proof for $b_{-}$ as the proof for $b_{+}$ follows similarly. By Lemma 2.8, it suffices to show for $i=1,2$ and $0\leq k\leq T_{1}-1$, each $S_{L_{-},\rho+\frac{\varepsilon}{2},\rho^{\prime}+\frac{\varepsilon}{2}}$-seminorm of $b_{i}\circ A^{k}$ is bounded uniformly in $k$ and $N$.
@@ 5.2.1. Porosity
As we shift by $z$, it suffices to only consider $I$ centered at the origin. Suppose towards a contradiction that $R$ and $\nu_{0}$ do not exist. Then there exists a sequence of points $z_{m}\in\mathbb{T}^{2n}$ and lines $I_{m}$ of length $m$ centered at the origin such that for all $\tau\in I_{m}$, $K_{j}$ does not contain $\{\varphi^{t}_{+}(\varphi^{\tau}_{+}(z_{m})):|t|\leq 1/m\}$. We pass to convergent subsequences and assume $z_{m}\rightarrow z_{\infty}$ and $I_{m}$ converges to $\mathbb{R}\tau$ for some $\tau\in\mathbb{R}^{l}$, i.e, the direction of $I_{m}$ converges to the direction of $\mathbb{R}\tau$. Rescale $\tau$ such that for $v_{\infty}=\sum_{i=1}^{l}\tau_{i}v_{i}^{+}$, $|v_{\infty}|=1$. As $K_{j}^{\circ}$ is safe, there exists $c\in\mathbb{R}$ such that $z_{\infty}+cv_{\infty}\mod\mathbb{Z}^{2n}\in K^{\circ}_{j}$. In other words, $\varphi_{+}^{c\tau}(z_{\infty})\in K_{j}^{\circ}$. Thus, for sufficiently large $m$, there exists $\tau_{m}\in I_{m}$ such that $\{\varphi^{t}_{+}(\varphi^{\tau_{m}}_{+}(z_{m})):|t|\leq 1/m\}\subset K_{j}^{\circ}$, a contradiction.
@@
As the right-hand side converges to $\infty$ as $h\rightarrow 0$, by taking sufficiently small $h$, we make the assumption that $0<j<T_{1}$.
@@
For sufficiently small $\varepsilon$, the right-hand-side converges to 0. Using (5.12) and the fact that $K_{w_{j}^{-}}\cap\operatorname{supp}b_{w_{j}^{-}}=\emptyset$, we know $A^{j}(\varphi^{t}_{+}(z)+v)\notin\operatorname{supp}b_{w_{j}^{-}}$ for sufficiently small $h$. From the definition of $b_{-}$, it follows that $\varphi^{t}_{+}(z)+v\notin\operatorname{supp}b_{-}$. ∎
@@ 5.2.2. Application of fractal uncertainty principle
Although $P_{k}$ is defined differently from §4 in this section, the original proof of [dyatlov2021semiclassical]*Lemma 4.5 holds for both definitions. On the other hand, although the statement of Lemma 5.7 is the same as [dyatlov2021semiclassical]*Lemma 4.6, the key ingredient at the end of the proof is different. In [dyatlov2021semiclassical], the authors assume a spectral gap condition for $A$. As a result, their lemma follows from an application of the 1-dimensional fractal uncertainty principle. However, as we made no such assumption, our lemma necessitates the higher-dimensional fractal uncertainty principle.
@@ 6.1. Short Periods
(6.2) $M_{A}f(x)=\frac{e^{-\frac{i\pi}{4}}}{\sqrt{2\pi hb}}\int_{\mathbb{R}}e^{\frac{i}{h}\left(\frac{d}{2b}x^{2}-\frac{xy}{b}+\frac{a}{2b}y^{2}\right)}f(y)dy.$
@@
We give a brief summary of their argument. Each summand in (6.7) with $|t|\leq P/4-o(P)$ behaves “locally,” concentrating at the origin in the semiclassical limit. On the other hand, the summands with $P/4+o(P)\leq|t|\leq P/2$ behave “ergodically,” equidistributing on the torus in the semiclassical limit. Finally, the terms close to $\pm P/4$ are between local and ergodic, but contribute negligibly in the semiclassical limit. Thus, the sum over $|t|\leq P/4$ weakly converges to $\tfrac{1}{2}\delta(x_{1},\xi_{1})$, while the sum over $P/4<|t|\leq P/2$ weakly converges to $\tfrac{1}{2}dx_{1}d\xi_{1}$.
@@ 6.3. Preliminary calculations
(6.8) $\begin{split}M_{A}G_{h}(x)&=(\pi h)^{-\frac{1}{4}}\frac{e^{-\frac{i\pi}{4}}}{\sqrt{2\pi hb}}e^{\frac{i}{2h}\left(\frac{d}{b}x^{2}\right)}\int_{\mathbb{R}}e^{\frac{i}{2h}\left(y^{2}\left(\frac{a}{b}+i\right)-\frac{2xy}{b}\right)}dy\\ &=(\pi h)^{-\frac{1}{4}}\frac{e^{-\frac{i\pi}{4}}}{\sqrt{2\pi hb}}e^{\frac{i}{2h}x^{2}\left(\frac{d}{b}-\frac{1}{b(a+ib)}\right)}\int_{\mathbb{R}}e^{\frac{i}{2h}y^{2}\left(\frac{a}{b}+i\right)}dy\\ &=(\pi h)^{-\frac{1}{4}}\frac{e^{-\frac{i\pi}{4}}}{\sqrt{b-ai}}e^{\frac{i}{2h}x^{2}\left(\frac{c+id}{a+ib}\right)}.\end{split}$
@@
$\left\langle U_{\omega}G_{h},M_{A}G_{h}\right\rangle_{L^{2}}=\frac{\sqrt{2}}{\sqrt{\operatorname{Tr}(A)}}e^{-\frac{1}{2h\operatorname{Tr}(A)}\left\langle A^{-1}\omega,\omega\right\rangle}e^{\frac{i}{2h\operatorname{Tr}(A)}\left\langle AJ\omega,\omega\right\rangle},$
where $J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$.
@@
(6.10) $\begin{split}\left\langle U_{\omega}G_{h},M_{A}G_{h}\right\rangle_{L^{2}}&=e^{\frac{i\pi}{4}}\frac{(\pi h)^{-\frac{1}{2}}}{\sqrt{ai+b}}e^{-\frac{i}{2h}y\eta}\int_{\mathbb{R}}e^{\frac{i}{h}\eta x-\frac{(x-y)^{2}}{2h}}\overline{e^{\frac{i}{2h}x^{2}\left(\frac{c+id}{a+ib}\right)}}dx\\ &=e^{\frac{i\pi}{4}}\frac{(\pi h)^{-\frac{1}{2}}}{\sqrt{ai+b}}e^{-\frac{i}{2h}y\eta}e^{-\frac{y^{2}}{2h}}\int_{\mathbb{R}}e^{-\frac{x^{2}}{2h}\left(\frac{a+d}{a-ib}\right)+\frac{x}{h}(i\eta+y)}dx\\ &=e^{\frac{i\pi}{4}}\frac{(\pi h)^{-\frac{1}{2}}}{\sqrt{ai+b}}e^{-\frac{i}{2h}y\eta}e^{-\frac{y^{2}}{2h}}e^{\frac{(i\eta+y)^{2}}{2h}\frac{a-ib}{a+d}}\int_{\mathbb{R}}e^{-\frac{x^{2}}{2h}\frac{a+d}{a-ib}}dx\\ &=\frac{\sqrt{2}}{\sqrt{a+d}}e^{-\frac{i}{2h}y\eta}e^{-\frac{y^{2}}{2h}}e^{\frac{(i\eta+y)^{2}}{2h}\frac{a-ib}{a+d}}\\ &=\frac{\sqrt{2}}{\sqrt{a+d}}e^{-\frac{1}{2h(a+d)}\left(a\eta^{2}-2b\eta y+dy^{2}\right)}e^{\frac{i}{2h(a+d)}(b\eta^{2}-by^{2}+(a-d)\eta y)}\\ &=\frac{\sqrt{2}}{\sqrt{\operatorname{Tr}(A)}}e^{-\frac{1}{2h\operatorname{Tr}(A)}\left\langle A^{-1}\omega,\omega\right\rangle}e^{\frac{i}{2h\operatorname{Tr}(A)}\left\langle AJ\omega,\omega\right\rangle},\end{split}$
for $J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$. ∎
@@
There exists $0<c_{0}<1$, depending only on $A$, such that for all $l\in\mathbb{Z}^{2}\setminus 0$,
@@
Using the above lemma, we can prove the following. The result generalizes [FNdB]*Proposition 1 to allow for a perturbation $r$.
@@
Suppose $r\in\mathbb{R}^{2}$. Then for all $q\in\mathbb{Z}$,
(6.13) $\sum_{l\in\mathbb{Z}^{2}}\left|\left\langle M^{q}_{A}G_{h},U_{l+r}G_{h}\right\rangle_{L^{2}}\right|\leq C(h\lambda^{|q|/2}+\lambda^{-|q|/2}).$
For $|r|\ll\min(1,\lambda^{-|q|}(h\log h^{-1})^{-\frac{1}{2}})$ and $q\in\mathbb{Z}$ with $h\log h^{-1}\ll\lambda^{-|q|}$,
(6.14) $\sum_{l\in\mathbb{Z}^{2}\setminus 0}\left|\left\langle M^{q}_{A}G_{h},U_{l+r}G_{h}\right\rangle_{L^{2}}\right|\leq Ch\lambda^{|q|/2}.$
Finally, for $|r|\ll 1$ and all $q\in\mathbb{Z}$,
(6.15) $\sum_{l\in\mathbb{Z}^{2}\setminus 0}\left|\left\langle M^{q}_{A}G_{h},U_{l+r}G_{h}\right\rangle_{L^{2}}\right|\leq C(h\lambda^{|q|/2}+(h\log h^{-1})^{\frac{1}{4}}).$
For each inequality, $C$ does not depend on $r$ or $q$.
@@
$\left|\left\langle M_{A}^{q}G_{h},U_{l+r}G_{h}\right\rangle_{L^{2}}\right|=\frac{\sqrt{2}}{\sqrt{\lambda^{q}+\lambda^{-q}}}e^{-\frac{a_{+}(l+r)^{2}\lambda^{-q}+a_{-}(l+r)^{2}\lambda^{q}}{2h(\lambda^{q}+\lambda^{-q})}}.$
We assume that $q\geq 0$. If instead $q<0$, the following argument is the same, up to exchanging $a_{+}(\cdot)$ and $a_{-}(\cdot)$ and replacing $q$ with q-q.
@@
$\left\{z\in\mathbb{R}^{2}:k-1\leq\frac{a_{+}(z+r)^{2}\lambda^{-q}+a_{-}(z+r)^{2}\lambda^{q}}{2h(\lambda^{q}+\lambda^{-q})}<k\right\}$
$\quad\subset\left\{z\in\mathbb{R}^{2}:\lambda^{-2q}a_{+}(z+r)^{2}+a_{-}(z+r)^{2}<4kh\right\}$
$\quad\subset\left\{z\in\mathbb{R}^{2}:|a_{+}(z+r)|<2\lambda^{q}\sqrt{kh}\text{ and }|a_{-}(z+r)|<2\sqrt{kh}\right\}\eqqcolon S_{k}.$
@@
(6.16) $\sum_{l\in\mathbb{Z}^{2}}\left|\left\langle M^{q}_{A}G_{h},U_{l+r}G_{h}\right\rangle_{L^{2}}\right|\leq C\lambda^{-q/2}\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}.$